OpenStudy (hiyathere_20):

Solve the quadratic equation x2 + 16x = -64 by factoring.

OpenStudy (hiyathere_20):

a. {2, 4} b. 8 c. {-8, 8} d. -8

OpenStudy (mathmale):

Correct: x^2 Wrong: x2

OpenStudy (mathmale):

Solve by factoring: x^2 + 16x = -64

OpenStudy (hiyathere_20):

yes that is correct.

OpenStudy (danjs):

For x^2 + 16x + 64 = 0 Try to think, what factors of 64 can be added or subtracted to give you the 16. 64 1 32 2 16 4 8 8 ...

OpenStudy (hiyathere_20):

4, 14

OpenStudy (danjs):

4 and 14 arent factors of 64, you are trying to make this look like (x + a)*(x + b) = x^2 + (a+b)*x + (a*b) = 0 where a and b are factors of 64 and total 16. a*b=64 a + b =16

OpenStudy (danjs):

Some factors of 64 are, 64 1 32 2 16 4 8 8 notice 8 and 8 total up to the 16 needed 8 * 8 = 64 8 + 8 = 16 x^2 + 16x + 64 = 0 (x + 8)*(x + 8) = 0

OpenStudy (mathmale):

Please involve yourself here. Name some of the methods of solution of quadratic equations that you know, and try applying one or two of them to this particularly problem. Then I'd be happy to give you feedback on your work.